OpenStudy (ikou):

What are the vertical asymptotes of the function f(x) = the quantity of 2 x plus 8, all over x squared plus 5 x plus 6?

OpenStudy (ikou):

2x+8 / x^2+5x+6

OpenStudy (ikou):

The answer choices x = −3 and x = −2 x = −3 and x = 2 x = 1 and x = −2 x = 1 and x = 2

OpenStudy (ikou):

I think its x = 1 and x = −2

OpenStudy (fortytherapper):

Mmm, it's not your choice. Would you know how to go about finding them?

OpenStudy (anonymous):

To find the vertical asymptotes, set the denominator equal to zero and solve. \[x ^{2}+5x+6=(x+3)(x+2)\] x=-2, x=-3

OpenStudy (anonymous):

That equation should be equal to zero, my mistake.

OpenStudy (witty_roger):

x=-2,-3 are the poles/asymptotes of the function

OpenStudy (ikou):

Oh that makes sense

OpenStudy (ikou):

so would that be the answer?

OpenStudy (witty_roger):

a

OpenStudy (ikou):

mkay ty guys